Q1. Suppose that the residents of Vegopia spend all of their income on long beans, bean
sprouts, and carrots. In 2006, they buy 100 strands of long beans for RM200.00, 50 kilos of
beans sprouts for RM75.00, and 500 carrots for RM50.00. In 2007, they buy 75 strands of
long beans for RM225.00, 80 kilos of beans sprouts for RM120.00, and 500 carrots for
RM100.00.
a. Calculate the price of each vegetable in each year.
b. Using 2006 as the base year, calculate the CPI for each year.
c. What is the inflation rate in 2007?
1
Expert's answer
2020-12-08T09:56:04-0500

(a) In 2006

Price of long beans "= \frac{200}{100} = 2 \; per \; strands"

Price of bean sprouts "= \frac{75}{50} = 1.5 \;per \;kg"

Price of carrots "= \frac{50}{500} = 0.1 \; per \; carrot"

In 2007

Price of long beans "= \frac{225}{75} = 3 \;per \;strands"

Price of bean sprouts "= \frac{120}{80} = 1.5 \;per \;kg"

Price of carrots "= \frac{100}{500} = 0.2 \;per \;carrot"

(b) As 2006 as base year,

The CPI of 2006 = 100

CPI in 2007 "= \frac{value \; of \; basket \; in \; 2007}{Value \; of \; basket \; in \; 2006} \times100"

The consumption quantities of 2006, will be considered in 2007.

CPI in 2007 "= \frac{(100 \times 3)+(50 \times 1.5)+(500 \times 0.2)}{(100 \times 2)+(50 \times 1.5)+(500\times 0.1)} \times100"

CPI in 2007 = 146.15

(c) Inflation rate in 2007 "= \frac{146.15 – 100}{100} = 46.15 \;persent"

It is the % increase the basket of goods in 2007, in comparison to the price of the same basket in 2006.